OpenStudy (anonymous):

x/x-1 +1 divided by x+2/x you have to get a common denominator in the numberator and muliply by the reciprocal

OpenStudy (anonymous):

\[\frac{\frac{x}{x-1}+1}{x+\frac{2}{x}}\]

OpenStudy (anonymous):

you can clear the fractions in one step by multiplying top and bottom by \[x(x-1)\]

OpenStudy (anonymous):

you get \[\frac{x^2+x(x-1)}{x^2(x-1)+2(x-1)}\] and then multiply out and combine like terms in numerator and denominator

OpenStudy (anonymous):

giving \[\frac{2x^2-x}{x^3-x^2+2x-2}\]